Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?
Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3050)

1/.5+1/.75+1/1.5=1/X
(3+2+1)/(.5*.75*1.5)=1/X
6/.5625=1/X
6X=.5625
X=.5625/6
X=.094 HOURS OR 5.64 MINUTES.

Answer by Alan3354(30993)

Jim does 1/30 of the job per minute
Sue does 1/45
Tony does 1/90
--------
Together, they do 1/30 + 1/45 + 1/90 = 6/90 per minute = 1/15
--> Together, it takes them 15 minutes.